A store brought a $1000 shipment of sweaters and offered them for sale after a $15 markup on each sweater. All but 10 sweaters were sold, and the store made a $400 profit on the sweaters. How many sweaters did the store buy?
Answer by stanbon(57400)

A store brought a $1000 shipment of sweaters and offered them for sale after a $15 markup on each sweater. All but 10 sweaters were sold, and the store made a $400 profit on the sweaters. How many sweaters did the store buy?
----
Let the # of sweaters bought be "x".
---
Average price paid was 1000/x dollars
-----
Markup price was (1000/x)+15 dollars
-------
Profit equation:
value of those sold - cost = profit
-----
(x-10)[(1000/x)+15] - 1000 = 400
-------
(x-10)[(1000+15x)/x] = 1400
---
(1-(10/x)(1000+15x) = 1400
----
x = 25 (# of shirts purchased)
===============================
Cheers,
Stan H.